OpenStudy (anonymous):

A student needs to make a square cardboard piece. The cardboard should have a perimeter equal to at least 92 inches. The function f(s) relates the perimeter of a cardboard piece, in inches, to the length of its side in inches. Which of the following shows a reasonable domain for f(s)? 23 < s < 46 23 < s < 92 s ≥ 92 s ≥ 23

OpenStudy (anonymous):

23 < s < 46 23 < s < 92 s ≥ 92 s ≥ 23

OpenStudy (anonymous):

you want the perimeter (p) to be greater than 92. the perimeter is the sum of the 4 sides (s), which on a square are all the same. so, p =4s which could be rearranged as s= p/4 so dividing the given minimum perimeter (92) by 4 will give you the smallest allowable side. so, s≥p/4

OpenStudy (anonymous):

so its c

OpenStudy (xapproachesinfinity):

f(s)>=92 f(s)=4s as we now from the perimeter of a square so s>=92/4

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

your perimeter is 92. or more your side is 1/4th of the perimeter so your minimum side must be = 92/4 and since the perimeter is allowed to be larger than 92, the side can also be > 92/4 find 92 /4, that is s s>=92/4

OpenStudy (anonymous):

ohh